OpenStudy (anonymous):

Find x if f(x) = 2x + 7 and f(x) =-1

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

Replace f(x) with -1 f(x) = 2x + 7 -1 = 2x + 7

OpenStudy (anonymous):

jk, its -4 sorry.

jimthompson5910 (jim_thompson5910):

Now solve for x -1 = 2x + 7 -1-7 = 2x + 7-7 ... Subtract 7 from both sides. -8 = 2x 2x = -8 what's the next step?

OpenStudy (anonymous):

I don't know, I'm confused with this stuff

jimthompson5910 (jim_thompson5910):

2x means 2 times x

jimthompson5910 (jim_thompson5910):

when we have 2x = -8, we divide both sides by 2 to isolate x we divide to undo the multiplication of 2 to x

jimthompson5910 (jim_thompson5910):

2x = -8 2x/2 = -8/2 1x = -4 x = -4

OpenStudy (anonymous):

So it's -4?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

-1=2x+7 2x=-1-7=-8 x=-4

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)